# Tracking Any Object Amodally

### 19 Dec 2023 v1

@@ 1 Introduction
Machine perception, particularly in object detection and tracking, has focused primarily on reasoning about visible or modal objects. This modal perception ignores parts of the three-dimensional world that are occluded to the camera. However, amodal completion of objects in the real-world (e.g., seeing a setting sun but understanding it is whole) and their persistence over time (e.g., person walking behind a car in Fig. 1) are fundamental capabilities that develop in humans in their early years [27, 40, 3]. In autonomous systems, this online amodal reasoning finds a direct application in downstream motion planning and navigation. Despite this, object detection and tracking stacks give little importance to partially or completely occluded objects; this becomes apparent in datasets that are only annotated modally [54, 30, 21, 36, 15, 16, 31, 58, 10, 9] but are still widely used and built upon by algorithms. These algorithms in turn learn to perceive only modal objects.
@@ 3 Dataset Annotation
Since TAO tracks (and their individual bounding boxes) are modal in nature, extending boxes to account for in-frame and out-of-frame occlusions translates to (1) modifying existing boxes in TAO, for the cases of partial occlusion, and (2) adding new boxes for completely occluded objects. We find that out of a total of 358,862 boxes in TAO, 266,902 (74.4%) boxes are modified to incorporate partial occlusion. TAO-Amodal boasts of an additional 23,449 bounding boxes for invisible objects that were not in TAO. These annotations follow the guidelines detailed in the supplementary, that cover a wide range of both in-frame and out-of-frame occlusion scenarios. Importantly on a higher-level, we only consider the occlusion cases, where an object has appeared in the scene before. We also scope out the occlusion cases, where an object might be partially behind the camera, or outside the annotation workspace defined above. Within the strict purview of the guidelines, when an object’s location cannot be discerned confidently by the annotators, an is_uncertain flag is added to the labeled bounding box. From the 23,449 boxes for invisible objects, 20,218 (85.8%) boxes are annotated confidently (i.e., without the uncertain flag), indicating that there is inherent uncertainty in localizing objects when they undergo heavy occlusions (ref. human vision experiment from a prior work [29]).
@@ 4.2 Crafting occlusion with PasteNOcclude (PnO)
To simulate occlusion scenarios during training, we introduced a data augmentation technique called PasteNOcclude (PnO). PasteNOcclude functions by integrating segments of objects into the original images to act as occludees. The segment collection comprises 505k objects extracted from LVIS [21] and COCO [36] images using segmentation masks. For each input image, we randomly select 1 to 7 segments from the collection and paste them at arbitrary locations, allowing for partial extension beyond the image boundary to replicate out-of-frame situations. Subsequently, we incorporate the ground truth boxes of the pasted segments into the original ground truth sets and apply regression loss according to the previously described pipeline. We find that PnO leads to improvements improvements in detection across all occlusion scenarios, shown in Table 3. We posit that this synthetic strategy is particularly important for the long-tailed nature of TAO-amodal, unlike COCO-amodal, where a similar synthetic occlusion strategy leads to limited improvement [62]. We provide visual examples of such synthetic occlusions in the supplement.
@@ 5 Empirical analysis
Our experiments are categorized into two parts. In Section 5.1, we assess the challenges of amodal detection and tracking by evaluating various off-the-shelf trackers. The efficacy of the amodal expander is examined in Section 5.2. Implementation details and ablation study could be found in the supplementary material.
@@ 5.1 Benchmarking state-of-the-art trackers
Using the estimated visibility attributes, we assess the tracking and detection capabilities of the model through variations of detection AP [36] and Track-AP [10], representing the average precision across all categories at an IoU threshold of 0.5. We label objects with visibility less than 0.1 as heavily occluded, evaluated as AP[0.0, 0.1], where the superscript indicates the range of object visibility. If the visibility falls between 0.1 and 0.8, we categorize them as partially occluded, while those with visibility greater than 0.8 are considered non-occluded. Objects that stretch partially beyond the image boundary are referred to as out-of-frame (OoF) and evaluated with APOoF. Additionally, we assess the model’s performance on modal annotations with Modal AP. In tracking, we evaluate highly or partially occluded tracks (Track-AP[0, 0.8]), which are track with visibility at or below 0.8 for more than 5 frames (seconds). We also evaluate performance on modal annotations (Modal Track AP). A table is provided for easy reference to each metric definition in the supplementary.
@@
We report detection and tracking average precision (AP) numbers of SOTA off-the-shelf trackers on TAO-Amodal validation set running at 1fps with an IoU threshold 0.5 in Tab. 2. We also observed similar performance trends when running at 5fps with higher IoU thresholds, shown in the supplementary. Every off-the-shelf tracker was trained on either TAO [10] or LVIS [21], ensuring alignment of category vocabulary with our dataset. GTR [61] was trained on the combination of LVIS and COCO [36] by generating synthetic videos using the training strategy in [60]. ViTDet [35] was trained on LVIS and combined with online SORT [4] tracker. We reproduced AOA [12] using their released implementation, with object detector trained on LVIS and tracking ReID head trained on TAO. QDTrack [17] and TET [34] followed similar training procedures, pretraining detectors on LVIS and training instance similarity heads on TAO.
@@
Existing off-the-shelf trackers exhibit reasonable performance in detecting non-occluded objects, with ViTDet achieving 39.71% AP[0.8,1] as revealed in Table 2. However, all trackers face challenges in detecting heavily occluded objects, and there is a noticeable decline in performance in partially occluded (AP[0.1,0.8]) and out-of-frame (OoF) scenarios. We noticed that ViTDet operating at 5 fps benefits from the property of SORT to estimate the location in the current frame using past information, resulting in improved results shown in the supplementary. Nevertheless, this improvement comes at the cost of processing ViT-Det on 5x more frames than models running at 1 fps. In contrast, amodal completion could be a promising way for efficiently handling occlusion.
@@ 5.2 Amodal expander experiments
Amodal expander architecture could be referred in Fig. 3. We built the expander on top of GTR [61] as this method shows reasonable performance in both detection and tracking aspects in Table 2. The hidden dimension of MLP is selected as 256. ReLU [1] and dropout [45] with a probability of 0.2 were applied to each layer except the last one. The training of the amodal expander was conducted on the TAO-Amodal training set, along with PasteNOcclude (PnO) and augmentation used in [61]. All the modules except the amodal expander are frozen during training. Other ablation study, hyperparameter details for training and PnO could be found in the supplementary material.
@@ 6 Discussion
In this supplement, we extend our discussion of the proposed dataset and method within the context of tracking any object with amodal perception. Specifically, we discuss details about the training and PasteNOcclude technique in Section 7, provide further empirical analysis in Section 8, and outline the annotation guidelines of our dataset in Section 9. We further provide comprehensive video demonstration of our dataset and qualitative results in webpage/index.html.

### 23 Jan 2024 v2

@@ 1 Introduction
Machine perception, particularly in object detection and tracking, has focused primarily on reasoning about visible or modal objects. This modal perception ignores parts of the three-dimensional world that are occluded to the camera. However, amodal completion of objects in the real-world (e.g., seeing a setting sun but understanding it is whole) and their persistence over time (e.g., person walking behind a car in Fig. 1) are fundamental capabilities that develop in humans in their early years [28, 43, 3]. In autonomous systems, this online amodal reasoning finds a direct application in downstream motion planning and navigation. Despite this, object detection and tracking stacks give little importance to partially or completely occluded objects; this becomes apparent in datasets that are only annotated modally [57, 31, 21, 38, 15, 16, 32, 61, 10, 9] but are still widely used and built upon by algorithms. These algorithms [37, 17, 64, 12, 41, 34, 25, 36, 46] in turn learn to perceive only modal objects.
@@ 3 Dataset Annotation
Since TAO tracks (and their individual bounding boxes) are modal in nature, extending boxes to account for in-frame and out-of-frame occlusions translates to (1) modifying existing boxes in TAO, for the cases of partial occlusion, and (2) adding new boxes for completely occluded objects. We find that out of a total of 358,862 boxes in TAO, 266,902 (74.4%) boxes are modified to incorporate partial occlusion. TAO-Amodal boasts of an additional 23,449 bounding boxes for invisible objects that were not in TAO. These annotations follow the guidelines detailed in the appendix, that cover a wide range of both in-frame and out-of-frame occlusion scenarios. Importantly on a higher-level, we only consider the occlusion cases, where an object has appeared in the scene before. We also scope out the occlusion cases, where an object might be partially behind the camera, or outside the annotation workspace defined above. Within the strict purview of the guidelines, when an object’s location cannot be discerned confidently by the annotators, an is_uncertain flag is added to the labeled bounding box. From the 23,449 boxes for invisible objects, 20,218 (85.8%) boxes are annotated confidently (i.e., without the uncertain flag), indicating that there is inherent uncertainty in localizing objects when they undergo heavy occlusions (ref. human vision experiment from a prior work [30]).
@@ 4.2 Crafting occlusion with PasteNOcclude (PnO)
To simulate occlusion scenarios during training, we introduced a data augmentation technique called PasteNOcclude (PnO). PasteNOcclude functions by integrating segments of objects into the original images to act as occludees. The segment collection comprises 505k objects extracted from LVIS [21] and COCO [38] images using segmentation masks. For each input image, we randomly select 1 to 7 segments from the collection and paste them at arbitrary locations, allowing for partial extension beyond the image boundary to replicate out-of-frame situations. Subsequently, we incorporate the ground truth boxes of the pasted segments into the original ground truth sets and apply regression loss according to the previously described pipeline. We find that PnO leads to improvements improvements in detection across all occlusion scenarios, shown in Table 3. We posit that this synthetic strategy is particularly important for the long-tailed nature of TAO-amodal, unlike COCO-amodal, where a similar synthetic occlusion strategy leads to limited improvement [65]. We provide visual examples of such synthetic occlusions in the appendix.
@@ 5 Empirical analysis
Our experiments are categorized into two parts. In Section 5.1, we assess the challenges of amodal detection and tracking by evaluating various off-the-shelf trackers. The efficacy of the amodal expander is examined in Section 5.2. Implementation details and ablation study could be found in the appendix.
@@ 5.1 Benchmarking state-of-the-art trackers
Using the estimated visibility attributes, we assess the tracking and detection capabilities of the model through variations of detection AP [38] and Track-AP [10], representing the average precision across all categories at an IoU threshold of 0.5. We label objects with visibility less than 0.1 as heavily occluded, evaluated as AP[0.0, 0.1], where the superscript indicates the range of object visibility. If the visibility falls between 0.1 and 0.8, we categorize them as partially occluded, while those with visibility greater than 0.8 are considered non-occluded. Objects that stretch partially beyond the image boundary are referred to as out-of-frame (OoF) and evaluated with APOoF. Additionally, we assess the model’s performance on modal annotations with Modal AP. In tracking, we evaluate highly or partially occluded tracks (Track-AP[0, 0.8]), which are track with visibility at or below 0.8 for more than 5 frames (seconds). We also evaluate performance on modal annotations (Modal Track AP). A table is provided for easy reference to each metric definition in the appendix.
@@
We report detection and tracking average precision (AP) numbers of SOTA off-the-shelf trackers on TAO-Amodal validation set running at 1fps with an IoU threshold 0.5 in Tab. 2. We also observed similar performance trends when running at 5fps with higher IoU thresholds, shown in the appendix. Every off-the-shelf tracker was trained on either TAO [10] or LVIS [21], ensuring alignment of category vocabulary with our dataset. GTR [64] was trained on the combination of LVIS and COCO [38] by generating synthetic videos using the training strategy in [63]. ViTDet [37] was trained on LVIS and combined with online SORT [4] tracker. We reproduced AOA [12] using their released implementation, with object detector trained on LVIS and tracking ReID head trained on TAO. QDTrack [17] and TET [36] followed similar training procedures, pretraining detectors on LVIS and training instance similarity heads on TAO.
@@
Existing off-the-shelf trackers exhibit reasonable performance in detecting non-occluded objects, with ViTDet achieving 39.71% AP[0.8,1] as revealed in Table 2. However, all trackers face challenges in detecting heavily occluded objects, and there is a noticeable decline in performance in partially occluded (AP[0.1,0.8]) and out-of-frame (OoF) scenarios. We noticed that ViTDet operating at 5 fps benefits from the property of SORT to estimate the location in the current frame using past information, resulting in improved results shown in the appendix. Nevertheless, this improvement comes at the cost of processing ViT-Det on 5x more frames than models running at 1 fps. In contrast, amodal completion could be a promising way for efficiently handling occlusion.
@@ 5.2 Amodal expander experiments
Amodal expander architecture could be referred in Fig. 3. We built the expander on top of GTR [64] as this method shows reasonable performance in both detection and tracking aspects in Table 2. The hidden dimension of MLP is selected as 256. ReLU [1] and dropout [48] with a probability of 0.2 were applied to each layer except the last one. The training of the amodal expander was conducted on the TAO-Amodal training set, along with PasteNOcclude (PnO) and augmentation used in [64]. All the modules except the amodal expander are frozen during training. Other ablation study, hyperparameter details for training and PnO could be found in the appendix.
@@ 6 Discussion
In this appendix, we extend our discussion of the proposed dataset and method within the context of tracking any object with amodal perception. Specifically, we discuss details about the training and PasteNOcclude technique in Section 7, provide further empirical analysis in Section 8, and outline the annotation guidelines of our dataset in Section 9. We further provide comprehensive video demonstration of our dataset and qualitative results at https://tao-amodal.github.io.